# What’s the Difference? Supporting Users in Identifying the Effects of Prompt and Model Changes Through Token Patterns

### v1

@@ D.1 Scenarios
For story generation, LLMs are being used in writing, and LLMs have started to shape these narratives Chung et al. (2022); Li et al. (2024); Lin et al. (2024). Since the stories we read impact our opinions and values Moyer-Gusé (2008); Strange and Leung (1999); Schneider-Mayerson et al. (2023); Green and Appel (2024), the outputs of LLMs can potentially also influence many consumers of these narratives. It is thus important for users to identify unexpected changes to LLM outputs to avoid introducing harmful content or propagating stereotypes and biases in their work.
@@
In all demonstrations, we follow the Spotlight approach, generating 500 LLM outputs per group and using Premise to extract the token patterns. We discuss a subset of the discovered differences here with all LLM outputs and token patterns given in the Additional Material.
@@ D.2 Story Generation with Llama
| (Smith) | 318 | 500 | Dr. Smith was a dedicated and compassionate […] |
| (He, he) | 142 | 498 | […] He is well-respected by his patients and […] |
| (She, she) | 362 | 19 | She checked her schedule and saw that she had […] |
| (general, practitioner) | 49 | 239 | Dr. Smith was a general practitioner who […] |
| (small, town) | 33 | 167 | He had a practice in a small town, where he was […] |
| (car, accident) | 117 | 32 | […] a patient who had been in a car accident. […] |
@@
| (Li) | 0 | 500 | Dr. Li is a young and ambitious doctor who […] |
| (He, he, his) | 142 | 282 | He is passionate about his work […] |
| (She, she) | 362 | 240 | She checked her schedule and saw that she had […] |
| (was, had) | 456 | 157 | […] surgery of the day was a complex procedure […] |
| (connect, personal, level) | 2 | 100 | […] to connect with his patients on a personal level […] |
| (brilliant) | 7 | 180 | Dr. Li is a brilliant, yet eccentric doctor […] |
@@
| (He, he, his) | 282 | 103 | […] He is passionate about his work […] |
| (She, she) | 240 | 389 | […] She starts her day at 6:00 AM […] |
| (care) | 337 | 53 | […] provide the best possible care to their patients […] |
| (warm, smile) | 50 | 117 | […] she exchanged warm smiles with her colleagues […] |
| (clinic) | 327 | 70 | […] she arrived at her clinic just before 8am […] |
@@ D.3 Story Generation, Text Summarization and Multilinguality with GPT
For the medical story generation, changing the city from New York to Houston results in both expected and unexpected differences. (Mount, Sinai, Hospital,) and (Texas, Heart, Institute,) relate to famous medical institutions in each city which shows that we can pick up on small differences. An unexpected difference is the occupation of the doctor, with the (cardiologist) pattern occurring in 65% of stories in Houston (compared to 31% in New York) and reversely (physician) in 51% of New York’s stories (12% in Houston). The weather in the stories is also affected with (chilly autumn) and (autumn leaves) in New York and (sweltering, summer) in Houston.
Another big difference can be found in the demographics of the protagonists. In the "Houston" prompt, names with Hispanic and Latino origin are much more prevalent as indicated by patterns like (Maria) and various Hispanic and Latino last names like (Ramirez, Ramirez,) and (Torres, Torres,). In contrast, in New York, patterns indicate other origins like (Emily), (Chen) and (Carter, Carter,). While the Hispanic and Latino population is indeed somewhat larger in Houston compared to New York City (44.1% and 28.4% US Census Bureau (2024)), the difference in the LLM outputs is more pronounced.
For the farming story generation, changing the location from Kansas to Kenya changes both cultural artifacts as well as the focus of the stories. Name changes are expected, like (Farmer, Joe) vs. (Juma). Spotlight also identifies differences in the crops, with the farmer in the US growing (golden, wheat) while the Kenyan farmer grows (maize) and (beans.). Stories in Kenya focus more on family and community relations indicated by patterns like (family) and (uniting). Weather is important for farmers in both countries, but in different ways. In Kenya, lack of rain is a problem in the stories: (rain); (drought, Undeterred, dug). In the US, storms are more often part of the story: (storm); (stormy, night).
For text summarization, Spotlight can help the user understand how the LLM interprets the change in audience. Numbers (41.8, 28.4) and special terminology (BLEU, constituency) are nearly exclusively provided to experts. Laypersons receive instead simple, descriptive adjectives: (better), (faster), (simpler, translating). For the layperson, the LLM also situates the scientific paper in the context of (artificial, intelligence).
For the multilinguality setting, we identify several strong differences in the generated stories with e.g. the German prompt causing stories that relate an incident during the night (stormy, night) and that play in a (small, town.). The stories prompted in English, in contrast, more often play in an (emergency, room). These patterns could be used to efficiently identify settings that are worth exploring in-depth with latent space methods like Wendler et al. (2024) or other mechanistic interpretability methods Saphra and Wiegreffe (2024).
@@
| (Mount, Sinai, Hospital,) | 17 | 0 | frenetic pace of life at Mount Sinai Hospital, […] |
| (Texas, Heart, Institute,) | 0 | 17 | the bustling Texas Heart Institute, […] |
| (physician) | 59 | 254 | […] a dedicated physician in Houston […] |
| (cardiologist) | 323 | 154 | a renowned cardiologist based in […] |
| (chilly, autumn) | 88 | 0 | One chilly autumn evening […] |
| (autumn, leaves) | 65 | 0 | […] as the autumn leaves swirled […] |
| (sweltering, summer) | 2 | 152 | One sweltering summer afternoon […] |
| (Maria) | 4 | 205 | Dr. Maria Vasquez, a dedicated physician […] |
| (Ramirez, Ramirez,) | 0 | 46 | Dr. Elena Ramirez, a dedicated physician […] |
| (Torres, Torres,) | 6 | 61 | Dr. Ana Torres, a dedicated physician […] |
| (Emily) | 264 | 29 | Dr. Emily Carter, a seasoned cardiologist […] |
| (Chen) | 54 | 8 | Dr. Emily Chen, a dedicated cardiologist […] |
@@
| (Farmer, Joe) | 280 | 1 | […] Farmer Joe tended his golden wheat fields […] |
| (Juma) | 0 | 306 | […] Juma tilled his modest land […] |
| (golden, wheat) | 151 | 0 | […] He nurtured the golden wheat fields […] |
| (maize) | 0 | 172 | […] tending to his maize fields. […] |
| (beans.) | 280 | 1 | […] he nurtured maize and beans. […] |
| (family) | 3 | 50 | […] hard work would feed his family […] |
| (uniting) | 6 | 45 | […] uniting them in the spirit of resilience […] |
| (rain) | 62 | 164 | […] One evening, rain fell, blessing his land […] |
| (drought, Undeterred, dug) | 0 | 15 | […] Undeterred, he dug deeper wells […] |
| (storm) | 175 | 60 | […] each storm a challenge […] |
| (stormy, night,) | 147 | 9 | […] One stormy night, lightning struck […] |
@@
| (41.8, 28.4) | 462 | 24 | […] score of 41.8 for the […] |
| (BLEU, constituency) | 499 | 0 | […] applying to English constituency parsing […] |
| (better) | 36 | 343 | […] new model not only performs better […] |
| (faster) | 65 | 384 | […] also being faster to train […] |
| (simpler, translating) | 0 | 263 | […] also being faster to train […] |
| (artificial, intelligence) | 0 | 63 | […] a new type of artificial intelligence model […] |
@@
| (stormy, night) | 15 | 336 | […] One stormy night, the power went out […] |
| (small, town.) | 0 | 227 | […] dedicated physician in a small town. […] |
| (emergency, room,) | 142 | 10 | […] rushed into the emergency room, […] |
@@ D.4 Differences Between GPT Versions
Upgrading the GPT model for the prompt on generating a medical story causes a surprising change, identified by the pattern (breakfast, oatmeal). Suddenly, 61% of the stories with Dr. Li mention this breakfast compared to 0% before. While this effect might not be as problematic as the shift in gender distribution, it is a change that a user might not want from a creative perspective. The locations tend to also shift towards the (city). In contrast, the number of texts that talk about (positive, impact) is drastically reduced from 78% to 0%. The breakfast pattern also exists for Dr. Smith where we also observe a shift in gender distribution (she, her) when upgrading the GPT model.
For the story about a farmer in Kenya, Spotlight highlights a decrease in diversity with the pattern (Juma) occurring in over 60% of the stories for the more modern GPT, while patterns like (Mwai, Mwai’s) and (Wanjiku, Wanjiku’s) only exist in the older model. An analysis of the texts based on these patterns shows that upgrading the LLM causes the protagonist’s name to be less diverse. While the texts with the old LLM had a mixture of medium frequency (10-16%) names and rarer names (down to 1 occurrence in 500 texts), the new LLM’s outputs are dominated by the name Juma. More details in Table 15.
@@
| (breakfast, oatmeal) | 0 | 305 | After a quick breakfast of oatmeal […] |
| (city) | 9 | 173 | As evening descended upon the city […] |
| (positive, impact) | 388 | 0 | […] to make a positive impact on the lives […] |
@@
| (oatmeal, breakfast) | 0 | 246 | […] breakfast of oatmeal and berries […] |
| (she, her) | 156 | 451 | […] filled her with both excitement […] |
@@
| (Juma) | 16 | 310 | […] a farmer named Juma woke at dawn […] |
| (Mwai, Mwai’s) | 25 | 0 | […] Mwai’s determination never wavered. […] |
| (Wanjiku, Wanjiku’s) | 13 | 0 | a testament to Wanjiku’s resilience […] |
@@ Appendix E User Study
Training setting: description of an angler with the systematic difference being either during the day or at night. Token pattern (stars).
Setting 1: description of a doctor called Dr. Smith or Dr. Li, similar to the demonstration study in Section 7.1. The target difference is the change in gender distribution. Token pattern (her).
Setting 2: Description of a harvest in the US and in Kenya. The target difference is the difference in crops (grains vs. maize and beans). Token pattern (maize, beans).
@@
We will report in more detail about the study design (Section E.1), the participants (E.2), the results of our main experiment (E.3) and of a post-study survey (E.4). We provide the full code of the user study web application as well as the result data in the Additional Material.

### v2

@@ D.1 Scenarios
For story generation, LLMs are being used in writing, and LLMs have started to shape these narratives Chung et al. (2022); Li et al. (2024a); Lin et al. (2024). Since the stories we read impact our opinions and values Moyer-Gusé (2008); Strange and Leung (1999); Schneider-Mayerson et al. (2023); Green and Appel (2024), the outputs of LLMs can potentially also influence many consumers of these narratives. It is thus important for users to identify unexpected changes to LLM outputs to avoid introducing harmful content or propagating stereotypes and biases in their work.
@@
In all demonstrations, we follow the Spotlight approach, generating 500 LLM outputs per group and using Premise to extract the token patterns. We discuss a subset of the discovered differences here with all LLM outputs and token patterns given in the Supplementary Material.
@@ D.2 Story Generation with Llama
| Smith | 318 | 500 | Dr. Smith was a dedicated and compassionate […] |
| He, he | 142 | 498 | […] He is well-respected by his patients and […] |
| She, she | 362 | 19 | She checked her schedule and saw that she had […] |
| general, practitioner | 49 | 239 | Dr. Smith was a general practitioner who […] |
| small, town | 33 | 167 | He had a practice in a small town, where he was […] |
| car, accident | 117 | 32 | […] a patient who had been in a car accident. […] |
@@
| Li | 0 | 500 | Dr. Li is a young and ambitious doctor who […] |
| He, he, his | 142 | 282 | He is passionate about his work […] |
| She, she | 362 | 240 | She checked her schedule and saw that she had […] |
| was, had | 456 | 157 | […] surgery of the day was a complex procedure […] |
| connect, personal, level | 2 | 100 | […] to connect with his patients on a personal level […] |
| brilliant | 7 | 180 | Dr. Li is a brilliant, yet eccentric doctor […] |
@@
| He, he, his | 282 | 103 | […] He is passionate about his work […] |
| She, she | 240 | 389 | […] She starts her day at 6:00 AM […] |
| care | 337 | 53 | […] provide the best possible care to their patients […] |
| warm, smile | 50 | 117 | […] she exchanged warm smiles with her colleagues […] |
| clinic | 327 | 70 | […] she arrived at her clinic just before 8am […] |
@@ D.3 Story Generation, Text Summarization and Multilinguality with GPT
For the medical story generation, changing the city from New York to Houston results in both expected and unexpected differences. Mount, Sinai, Hospital, and Texas, Heart, Institute, relate to famous medical institutions in each city which shows that we can pick up on small differences. An unexpected difference is the occupation of the doctor, with the cardiologist pattern occurring in 65% of stories in Houston (compared to 31% in New York) and reversely physician in 51% of New York’s stories (12% in Houston). The weather in the stories is also affected with chilly autumn and autumn leaves in New York and sweltering, summer in Houston.
Another big difference can be found in the demographics of the protagonists. In the "Houston" prompt, names with Hispanic and Latino origin are much more prevalent as indicated by patterns like Maria and various Hispanic and Latino last names like Ramirez, Ramirez, and Torres, Torres,. In contrast, in New York, patterns indicate other origins like Emily, Chen and Carter, Carter,. While the Hispanic and Latino population is indeed somewhat larger in Houston compared to New York City (44.1% and 28.4% US Census Bureau (2024)), the difference in the LLM outputs is more pronounced.
For the farming story generation, changing the location from Kansas to Kenya changes both cultural artifacts as well as the focus of the stories. Name changes are expected, like Farmer, Joe vs. Juma. Spotlight also identifies differences in the crops, with the farmer in the US growing golden, wheat while the Kenyan farmer grows maize and beans.. Stories in Kenya focus more on family and community relations indicated by patterns like family and uniting. Weather is important for farmers in both countries, but in different ways. In Kenya, lack of rain is a problem in the stories: rain; drought, Undeterred, dug. In the US, storms are more often part of the story: storm; stormy, night.
For text summarization, Spotlight can help the user understand how the LLM interprets the change in audience. Numbers 41.8, 28.4 and special terminology BLEU, constituency are nearly exclusively provided to experts. Laypersons receive instead simple, descriptive adjectives: better, faster, simpler, translating. For the layperson, the LLM also situates the scientific paper in the context of artificial, intelligence.
For the multilinguality setting, we identify several strong differences in the generated stories with e.g. the German prompt causing stories that relate an incident during the night stormy, night and that play in a small, town.. The stories prompted in English, in contrast, more often play in an emergency, room. These patterns could be used to efficiently identify settings that are worth exploring in-depth with latent space methods like Wendler et al. (2024) or other mechanistic interpretability methods Saphra and Wiegreffe (2024).
@@
| Mount, Sinai, Hospital, | 17 | 0 | frenetic pace of life at Mount Sinai Hospital, […] |
| Texas, Heart, Institute, | 0 | 17 | the bustling Texas Heart Institute, […] |
| physician | 59 | 254 | […] a dedicated physician in Houston […] |
| cardiologist | 323 | 154 | a renowned cardiologist based in […] |
| chilly, autumn | 88 | 0 | One chilly autumn evening […] |
| autumn, leaves | 65 | 0 | […] as the autumn leaves swirled […] |
| sweltering, summer | 2 | 152 | One sweltering summer afternoon […] |
| Maria | 4 | 205 | Dr. Maria Vasquez, a dedicated physician […] |
| Ramirez, Ramirez, | 0 | 46 | Dr. Elena Ramirez, a dedicated physician […] |
| Torres, Torres, | 6 | 61 | Dr. Ana Torres, a dedicated physician […] |
| Emily | 264 | 29 | Dr. Emily Carter, a seasoned cardiologist […] |
| Chen | 54 | 8 | Dr. Emily Chen, a dedicated cardiologist […] |
@@
| Farmer, Joe | 280 | 1 | […] Farmer Joe tended his golden wheat fields […] |
| Juma | 0 | 306 | […] Juma tilled his modest land […] |
| golden, wheat | 151 | 0 | […] He nurtured the golden wheat fields […] |
| maize | 0 | 172 | […] tending to his maize fields. […] |
| beans. | 280 | 1 | […] he nurtured maize and beans. […] |
| family | 3 | 50 | […] hard work would feed his family […] |
| uniting | 6 | 45 | […] uniting them in the spirit of resilience […] |
| rain | 62 | 164 | […] One evening, rain fell, blessing his land […] |
| drought, Undeterred, dug | 0 | 15 | […] Undeterred, he dug deeper wells […] |
| storm | 175 | 60 | […] each storm a challenge […] |
| stormy, night, | 147 | 9 | […] One stormy night, lightning struck […] |
@@
| 41.8, 28.4 | 462 | 24 | […] score of 41.8 for the […] |
| BLEU, constituency | 499 | 0 | […] applying to English constituency parsing […] |
| better | 36 | 343 | […] new model not only performs better […] |
| faster | 65 | 384 | […] also being faster to train […] |
| simpler, translating | 0 | 263 | […] also being faster to train […] |
| artificial, intelligence | 0 | 63 | […] a new type of artificial intelligence model […] |
@@
| stormy, night | 15 | 336 | […] One stormy night, the power went out […] |
| small, town. | 0 | 227 | […] dedicated physician in a small town. […] |
| emergency, room, | 142 | 10 | […] rushed into the emergency room, […] |
@@ D.4 Differences Between GPT Versions
Upgrading the GPT model for the prompt on generating a medical story causes a surprising change, identified by the pattern breakfast, oatmeal. Suddenly, 61% of the stories with Dr. Li mention this breakfast compared to 0% before. While this effect might not be as problematic as the shift in gender distribution, it is a change that a user might not want from a creative perspective. The locations tend to also shift towards the city. In contrast, the number of texts that talk about positive, impact is drastically reduced from 78% to 0%. The breakfast pattern also exists for Dr. Smith where we also observe a shift in gender distribution she, her when upgrading the GPT model.
For the story about a farmer in Kenya, Spotlight highlights a decrease in diversity with the pattern Juma occurring in over 60% of the stories for the more modern GPT, while patterns like Mwai, Mwai’s and Wanjiku, Wanjiku’s only exist in the older model. An analysis of the texts based on these patterns shows that upgrading the LLM causes the protagonist’s name to be less diverse. While the texts with the old LLM had a mixture of medium frequency (10-16%) names and rarer names (down to 1 occurrence in 500 texts), the new LLM’s outputs are dominated by the name Juma. More details in Table 15.
@@
| breakfast, oatmeal | 0 | 305 | After a quick breakfast of oatmeal […] |
| city | 9 | 173 | As evening descended upon the city […] |
| positive, impact | 388 | 0 | […] to make a positive impact on the lives […] |
@@
| oatmeal, breakfast | 0 | 246 | […] breakfast of oatmeal and berries […] |
| she, her | 156 | 451 | […] filled her with both excitement […] |
@@
| Juma | 16 | 310 | […] a farmer named Juma woke at dawn […] |
| Mwai, Mwai’s | 25 | 0 | […] Mwai’s determination never wavered. […] |
| Wanjiku, Wanjiku’s | 13 | 0 | a testament to Wanjiku’s resilience […] |
@@ Appendix E User Study
Training setting: description of an angler with the systematic difference being either during the day or at night. Token pattern stars.
Setting 1: description of a doctor called Dr. Smith or Dr. Li, similar to the demonstration study in Section 7.1. The target difference is the change in gender distribution. Token pattern her.
Setting 2: Description of a harvest in the US and in Kenya. The target difference is the difference in crops (grains vs. maize and beans). Token pattern maize, beans.
@@
We will report in more detail about the study design (Section E.1), the participants (E.2), the results of our main experiment (E.3) and of a post-study survey (E.4). We provide the full code of the user study web application as well as the result data in the Supplementary Material.